# Combining Constrained Diffusion Models and Numerical Solvers for Efficient and Robust Non-Convex Trajectory Optimization

### v3

@@ 6 Conclusion and Discussion
This paper presents a general and parallelizable framework for non-convex trajectory optimization, using a novel constrained diffusion model combined with numerical solvers. It efficiently provides divesrse solutions for problems with mutiple local optima, with less constraint violation, and up to 30 times acceleration. It is a general acceleration framework for solving scientific optimization problems.
Limitations also exist for the current framework. Although most constraint functions in trajectory optimization are differentiable (almost everywhere), current CDM does not work for non-differentiable violation functions. For non-differentiable constraints, transforming them into differentiable functions or applying approximation with appropriate bump functions or neural networks is possible. Another problem is that the diffusion sampling may not be achieved in real time in a highly dynamic environment, which can be improved with stride sampling, DDIM [40], etc.
@@ A.1 Model Details
For both unconstrained and constrained diffusion models, we apply a UNet structure with three hidden layers of 512, 512 and 1024 neurons. We also use a fully connected layer of 256 and 512 neurons to embed the conditional input $y$. The sampling step is set to be 500. Both DMs are trained with 200 epochs using the Adam optimizer [26]. The constrained and unconstrained DMs usually take 14 hours and 6 hours respectively, as well as 20 GB VRAM to train on one NVIDIA A100 GPU, respectively.
A.2 Tabletop Manipulation
@@
In this task $\mathcal{P}_{y}$, we aim to minimize the time for the gripper to reach the goal while avoiding the obstacles. The start position is fixed at $(5,5)$. The conditional parameter $y$ includes the goal position $p_{\text{goal}}$ randomly chosen from $\{(1,1),(1,9),(9,1),(9,9)\}$, 4 obstacle positions $p_{\text{obs}}$ randomly sampled between the start and goal, and the obstacle radius $r_{\text{obs}}$ randomly sampled from $[0.5,1.0]$. We discretize this problem into $T$ timesteps where $T=80$. The variable to optimize is $x=(t_{\text{final}},u_{1},u_{2},..,u_{T})\in\mathbb{R}^{161}$, and the conditional parameter is $y=(p_{\text{goal}},p_{\text{obs}},r_{\text{obs}})\in\mathbb{R}^{14}$. The objective is defined as $J(x;y)=t_{\text{final}}$. The constraint $g$ includes the goal-reaching constraints at $t_{\text{final}}$ and the collision avoidance for obstacles at all time. We use the 4-th order Runge-Kutta (RK4) method to integrate the dynamics and formulate the constraints.
@@
In this task, to collect the training data we sample 2700 different $y$, each with 100 uniformly sampled initial guesses. Then we use the SNOPT [19] solver for each problem instance to collect locally optimal solutions. Finally, we filter out those solutions with objective values greater than 6 s and use the remaining 237k data as training data. Within this 237k data, 10% of the data are used as the validation set. The whole data collection process takes around 15 hours with 200 AMD EPYC 9654 CPU cores.
@@
[…]
Figure 7: An example of tabletop manipulation task for goal-reaching and obstacle avoidance movement, using a CDM with a numerical solver for planning: (left) starting position of the gripper and the planned trajectory; (right) target position reached by the gripper.
A.3 Two Car Reach-Avoid
Problem formulation and solving.
In this task $\mathcal{P}_{y}$, we aim to minimize the time for the two cars to reach each own goal while avoiding colliding with the other car and obstacles. We fixed the start $p_{\text{start}}$ and goal $p_{\text{goal}}$ position for car1 to be $(0,10)$ and $(10,0)$, and for car2 to be $(10,0)$ and $(0,0)$. The conditional parameters $y\in\mathbb{R}^{6}$ include the two obstacle positions $p_{\text{obs}}$ randomly sampled from $[2.0,8.0]$, and the obstacle radius $r_{\text{obs}}$ randomly sampled from $[0.5,1.5]$. We formulate this trajectory optimization problem with a forward shooting method, with the discretized time step $T=40$. The variable to optimize is $x=(t_{\text{final}},u^{1}_{1},u^{1}_{2},...,u^{1}_{T},u^{2}_{1},u^{2}_{2},...,u^{2}_{T})$, where $u^{1}_{i}$ and $u^{2}_{i}$ are controls for two cars, respectively. The problem parameter is $y=(p_{\text{obs}},r_{\text{obs}})$. The objective is defined as $J(x;y)=t_{\text{final}}$. The constraint $g$ includes the goal-reaching constraints at $t_{\text{final}}$, and the collision avoidance for obstacles and between each car for all time. We use the 4-th order Runge-Kutta (RK4) method to integrate the dynamics and formulate the constraints.
Data collection and model training.
In this task, we collect locally optimal solutions $x^{*}$ from 3k different $y$, each with 100 uniformly sampled initial guesses using SNOPT [19]. We filter out those solutions whose objective is higher than 12 s and use the remaining 114k solutions as the training data. Within this 114k data, 10% of the data are used as the validation set. The whole data collection process takes around 8 hours with 200 AMD EPYC 9654 CPU cores.
@@ A.4 Cislunar Transfer
$\bm{x}=(t_{\text{burn}},t_{\text{coast}}^{\text{initial}},t_{\text{coast}}^{\text{final}},m_{f},u_{1},u_{2},...,u_{N})$ (10)
where $t_{\text{burn}},t_{\text{coast}}^{\text{initial}},t_{\text{coast}}^{\text{final}}$ are the effective burning time for the engine, initial and final coast time, $m_{f}$ is the final mass, and $u_{i}$ is a three-dimensional thrust control of the spacecraft. The objective is defined as $J(x;y)=-m_{f}$, which is to minimize fuel expenditure. The constraint $g$ is the midpoint agreement when integrating dynamics from forward and backward. The problem parameter $y$ is the maximum allowable thrust for the spacecraft and is bounded by $[0.1,1]$ Newton. To integrate the dynamics, we use the Runge-Kutta-Fehlberg 5-4th order (RK54) method [17].
@@
To collect the training data, we sample 12 different $y$ values from a grid in $[0.1,1]$ Newton, and for each $y$ we collect 25k locally optimal solutions with $m_{f}\geq 415kg$ total amount of 300k data. The problem is solved using pydylan, a python interface of the Dynamically Leveraged Automated (N) Multibody Trajectory Optimization (DyLAN) [3] and the solver SNOPT [19]. Within this 300k data, 10% of the data are used as the validation set. The whole data collection process takes around 3 days with 500 Intel Skylake CPU cores.

### v4

@@ 6 Conclusion and Discussion
This paper presents DiffuSolve: a general diffusion model-based solver for non-convex trajectory optimization. It can generate locally optimal, feasible, and diverse solutions with high computational efficiency. We also propose DiffuSolve+ which includes a novel constrained diffusion model. It has fewer constraint violations as a standalone method and further improves the computational efficiency when integrated with DiffuSolve+. It is also a general acceleration framework that has the potential to solve optimization problems in other areas beyond trajectory optimization such as finance, computational chemistry, etc.
Limitations also exist for the current paper. We don’t spend extra effort to optimize the diffusion model’s sampling speed. As a result, the current DiffusSolve and CDM may not be able to achieve real-time optimization for high-dimensional systems or highly dynamic environments. This can be improved with existing acceleration techniques such as stride sampling, DDIM (Song et al., 2020a), etc. In addition, although most constraint functions in trajectory optimization are differentiable (almost everywhere), the current CDM does not work for non-differentiable violation functions. For non-differentiable constraints, transforming them into differentiable functions or applying approximation with appropriate bump functions or neural networks is possible.
@@ A.1 Model Details
For both vanilla and constrained diffusion models (DMs), we apply a UNet structure with three hidden layers of 512, 512, and 1024 neurons. We also use a fully connected layer of 256 and 512 neurons to embed the conditional input $y$. The sampling step is set to be 500. Both DMs are trained with 200 epochs using the Adam optimizer (Kingma & Ba, 2014). The constrained and unconstrained DMs usually take 14 hours and 9 hours respectively as well as 20 GB VRAM to train on one NVIDIA A100 GPU, for every 100k training data with batch size of 512.
A.2 Two Car Reach-Avoid
@@
In this task $\mathcal{P}_{y}$, we aim to minimize the time for the two cars to reach each own goal while avoiding colliding with the other car and obstacles. We fixed the start $p_{\text{start}}$ and goal $p_{\text{goal}}$ position for car1 to be $(0,10)$ and $(10,0)$, and for car2 to be $(10,0)$ and $(0,0)$. The conditional parameters $y\in\mathbb{R}^{6}$ include the two obstacle positions $p_{\text{obs}}$ randomly sampled from $[2.0,8.0]$, and the obstacle radius $r_{\text{obs}}$ randomly sampled from $[0.5,1.5]$. We formulate this trajectory optimization problem with a forward shooting method, with the discretized time step $T=20$. The variable to optimize for each car is $x=(t_{\text{final}},u^{1}_{1},u^{1}_{2},...,u^{1}_{T},u^{2}_{1},u^{2}_{2},...,u^{2}_{T})\in\mathbb{R}^{41}$, where $u^{1}_{i}$ and $u^{2}_{i}$ are two control variables for each cars. The problem parameter is $y=(p_{\text{obs}},r_{\text{obs}})$. The objective is defined as $J(x;y)=t_{\text{final}}$. The constraint $g$ includes the goal-reaching constraints at $t_{\text{final}}$, and the collision avoidance for obstacles and between each car for all time. We use the 4-th order Runge-Kutta (RK4) method to integrate the dynamics and formulate the constraints.
@@
In this task, we collect locally optimal solutions $x^{*}$ from 3k different $y$, each with 100 uniformly sampled initial guesses using SNOPT (Gill et al., 2005). We filter out those solutions whose objective is higher than 12 s and use the remaining 114k solutions as the training data. Within this 114k data, 10% of the data are used as the validation set. The whole data collection process takes around 8 hours with 200 AMD EPYC 9654 CPU cores.
A.3 Quadrotor Navigation
Problem formulation and solving.
In this task $\mathcal{P}_{y}$, we aim to minimize the time for the quadrotor to reach the goal while avoiding the obstacles. The start position is fixed at $(-12,0,0)$, and the reference goal position is $(12,0,0)$. The conditional parameter $y$ includes the perturbation $d_{\text{goal}}$ randomly chosen from $[-2,2]$ that is added to the goal position, 4 obstacle positions $p_{\text{obs}}$ randomly sampled between the start and goal, and the obstacle radius $r_{\text{obs}}$ randomly sampled from $[1.5,3.5]$. We discretize this problem into $T$ timesteps where $T=80$. The variable to optimize is $x=(t_{\text{final}},u_{1}^{1},..,u_{T}^{1},u_{1}^{2},...u_{T}^{2},u_{1}^{3},...u_{T}^{3})\in\mathbb{R}^{241}$, and the conditional parameter is $y=(d_{\text{goal}},p_{\text{obs}},r_{\text{obs}})\in\mathbb{R}^{17}$. The objective is defined as $J(x;y)=t_{\text{final}}$. The constraint $g$ includes the goal-reaching constraints at $t_{\text{final}}$ and the collision avoidance for obstacles at all time. We use the 4-th order Runge-Kutta (RK4) method to integrate the dynamics and formulate the constraints.
Data collection and model training.
In this task, to collect the training data we sample 4000 different $y$, each with 50 uniformly sampled initial guesses. Then we use the SNOPT (Gill et al., 2005) solver for each problem instance to collect locally optimal solutions. Finally, we filter out those solutions with objective values greater than 4.57 s and use the remaining 179k data as training data. Within this 179k data, 10% of the data are used as the validation set. The whole data collection process takes around 20 hours with 200 AMD EPYC 9654 CPU cores.
@@ A.4 Cislunar Transfer
$x=(t_{\text{burn}},t_{\text{coast}}^{\text{initial}},t_{\text{coast}}^{\text{final}},m_{f},u_{1},u_{2},...,u_{N})\in\mathbb{R}^{64}$ (11)
where $t_{\text{burn}},t_{\text{coast}}^{\text{initial}},t_{\text{coast}}^{\text{final}}$ are the effective burning time for the engine, initial and final coast time, $m_{f}$ is the final mass, and $u_{i}$ is a three-dimensional thrust control of the spacecraft. The objective is defined as $J(x;y)=-m_{f}$, which is to minimize fuel expenditure. The constraint $g$ is the midpoint agreement when integrating dynamics from forward and backward. The problem parameter $y$ is the maximum allowable thrust for the spacecraft and is bounded by $[0.1,1]$ Newton. To integrate the dynamics, we use the Runge-Kutta-Fehlberg 5-4th order (RK54) method (Fehlberg, 1969).
@@
To collect the training data, we sample 12 different $y$ values from a grid in $[0.1,1]$ Newton, and for each $y$ we collect 25k locally optimal solutions with $m_{f}\geq 415kg$ total amount of 300k data. The problem is solved using pydylan, a python interface of the Dynamically Leveraged Automated (N) Multibody Trajectory Optimization (DyLAN) (Beeson et al., 2022) and the solver SNOPT (Gill et al., 2005). Within this 300k data, 10% of the data are used as the validation set. The whole data collection process takes around 3 days with 500 Intel Skylake CPU cores.